The wheel of a bicycle is rotating from rest at a constant rate of 1.86 rad/s^2. How much is the increase in angular velocity (in rpm) every minute?


1
Expert's answer
2021-11-24T11:57:43-0500
"\omega=\omega_0+\alpha t=0+1.86\ \dfrac{rad}{s^2}\times60\ s=116.6\ \dfrac{rad}{s},""\omega=116.6\ \dfrac{rad}{s}\times\dfrac{60\ s}{1\ min}\times\dfrac{1\ rev}{2\pi\ rad}=1066\ \dfrac{rev}{min}."